OpenStudy (anonymous):

A uniform rod of mass M = 250.0 g and length L = 50.0 cm stands vertically on a horizontal table. It is released from rest to fall.

OpenStudy (anonymous):

Calculate the vertical acceleration of the moving end of the rod as it makes an angle θ = 45° with respect to the vertical (take upward as positive and downward as negative). Calculate the normal force exerted by the table on the rod as it makes an angle θ = 45° with respect to the vertical.

OpenStudy (anonymous):

my question is that is vertical acceleration constant? since the gravitational acc. is constant i thought vertical acc would be constant as well.

OpenStudy (anonymous):

I think what the question seeks is that you find moment of inertia of the rod, from its end, which is I=M L/12 + M (L/2)^2 and then calculate torque due to gravity acting on the center of mass, m g L sin(45), finally equating the torque with the product of I and the angular acceleration. Normal force on table seems to be m g sin(45) but am not sure, due to complexity motion.

OpenStudy (anonymous):

torque vector tau = g M (L/2) sin(45) = I alpha. alpha vector = angular acceleration = acceleration / (L/2) = a / (L/2) vector a has a horizontal component proportional to sin(45) and a vertical acceleration component proportional to - cos(45) Sorry it is hard to make these vectors and scalar clear here.

OpenStudy (anonymous):

steps makes good sense but i dont quite get how we find I of different shapes? I hope we dont solve the  integration everytime?